OpenStudy (anonymous):

by USING HALF ANGLE IDENTITIES,find the exact values of cos 7 pi/8

OpenStudy (anonymous):

cos 7 pi/8= Cos ( pie - pie/ 8) = -cos pie/8

OpenStudy (anonymous):

hiw did u get -cos?

OpenStudy (anonymous):

let 2x = pie/8 x = pie/4 cos2x = 2cos^2x - 1 - cos 2x = 1 - 2 cos^2 x

OpenStudy (anonymous):

Substitute the values

OpenStudy (anonymous):

(pie - pie/8) is in 2 nd quadrant

OpenStudy (anonymous):

that's how u get -cos?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but the method that u did is not how i am suppose to do i have to use the half angle identities can u show me how to do it with that identitis?

OpenStudy (anonymous):

PLZA

OpenStudy (raden):

im not more like to pi, actually i have to convert pi radian to degree, 7 pi/8= 157.5 deg = 315/2 deg :)

OpenStudy (raden):

after that use the half angles formula : cos(x/2) = sqrt[(1+cos(x))/2] here x=315, try it ...

OpenStudy (anonymous):

yeah that works but just tell show me how get right answer by using that identitie

OpenStudy (raden):

use the formula above ^ cos(315/2) = sqrt[(1+cos(315))/2] cos157.5 = sqrt[(1+1/2*sqrt(2))/2] = ... dont forget simplify :)

OpenStudy (anonymous):

k the answer my teacher got is